# Heisenberg spin chain models for realising quantum battery with the aid of Dzyaloshinskii–Moriya interaction

### v1

@@ IV.2.3 Supercube spin chain configuration
IV.2.4 Extending the battery to 12 qubits:
We extend the supercube architecture from 8 qubits to 12 qubits by adding an additional face to the original cube to investigate the effect of increasing the system size in this configuration. From the ergotropy plot shown in figure LABEL:fig-b:erg_12q, we observe that the system initially charges to a high-energy state which is slightly less than the ideal maximum of $2n\hbar\omega_{0}=24$. The subsequent discharging cycles show a decay in oscillation amplitude. Importantly, the 12-qubit system also relaxes to a state very close to the ground (passive) state after each cycle, leaving only a negligible amount of residual energy that cannot be extracted. This indicates that increasing the number of qubits does not disrupt the characteristic sinusoidal charging–discharging behavior. However, the cycle efficiency in the 12-qubit case is somewhat reduced compared to the 8-qubit supercube spin chain.
Beyond this supercube extension, we also explore alternative 12-qubit connectivities based on the cuboctahedron and icosahedron skeletons. The cuboctahedral structure yields an almost ideal sinusoidal curve. The cuboctahedral structure yields an almost ideal sinusoidal curve for $J=1$ and $D=1.94$, with minimal non-extractable residual energy, as illustrated in Fig. LABEL:fig-b:octa_erg. In the case of the icosahedron, we observe an even closer approximation to perfect charging–discharging cycles for $J=1$ and $D=2.06$, as shown in Fig. LABEL:fig-b:icosa_erg. Notably, this configuration also reaches the maximum ergotropy value more rapidly than the simple cubic extension.
V Conclusion
@@
In this work, we studied the charging performance of eight-qubit Heisenberg spin chain quantum batteries (QBs) under the influence of Dzyaloshinskii-Moriya interaction (DMI). We compared three distinct spin chain configurations, namely, open boundary, closed boundary, and supercube, under both Ising and XXZ Hamiltonians. Our results show that the supercube consistently exhibits superior performance among all three configurations. For the Ising model, DMI enhances ergotropy in both the open boundary and supercube spin chains. In the XXZ model, the supercube configuration achieves a stable, near-ideal sinusoidal charging–discharging cycle, enabling periodic charging with negligible residual energy per cycle for optimal interaction strengths $D=1.7$ and $J=1$. We also investigated the effect of including the battery Hamiltonian $H_{B}$ in the charging process through the influence parameter $\lambda$. For all configurations and models, increasing $\lambda$ reduces the peak ergotropy. Interestingly, in the XXZ model, ergotropy for $D=1.7$ and $\lambda>0$ is higher than the $\lambda=0$ case for most of the evolution.
Additionally, we explored the effect of structural modifications of the supercube configuration. Introducing extra interactions between qubits disrupts the oscillatory behavior observed in the XXZ model. Extending the system size to 12 qubits preserves the sinusoidal nature of the ergotropy curve, but the system no longer discharges fully to its ground state, resulting in residual energy stored in the system. Notably, the 12-qubit icosahedral structure revives a perfect charging–discharging cycle, fully returning to the ground state without residual energy, thereby recovering the efficiency lost in the simple cubic extension. Although entanglement often influences QB performance, it has been shown that entanglement generation is not necessary for optimal work extraction [60]. In line with this, although our study does not explicitly examine entanglement, the results demonstrate enhancements in ergotropy and power without any direct connection to entanglement. Our findings provide insights into designing scalable and efficient quantum batteries and guide future exploration of optimized interaction topologies and system sizes. Interestingly, supercube-configured spin chain quantum batteries achieve high-efficiency, near-ideal charging cycles without requiring entanglement, providing a scalable route for next-generation quantum energy storage.
@@
The used codes and data sets are available in the GitHub repository. The code provided utilizes the QuantumToolbox package in both Python and Julia. [61, 62].

### v2

@@ III.2.3 Supercube spin chain configuration
No. of qubits (n) No. of connections (c) DMI strength (D)
3 2 1.86
4 3 1.96
5 4 2.03
6 4 2.18
5 2.07
7 6 2.11
8 5 2.30
6 2.23
7 2.13
9 6 2.36
8 2.15
10 7 2.33
8 2.25
9 2.17
11 8 2.35
10 2.18
12 7 2.52
8 2.46
9 2.33
10 2.27
11 2.20
Table 1: Different quantum battery structures with their corresponding DMI strength (D) for which ergotropy follows a sinusoidal curve
III.2.4 Structures with different number of qubits and coordination number
While the 8-qubit supercube achieves a near-ideal sinusoidal charging-discharging cycle, it is important to assess whether this behavior persists when the system size and coordination numbers are varied. To this end, we examine a family of $c$-regular graphs with qubit numbers ranging from $n=3$ to $n=12$, where each qubit interacts with exactly $c$ neighbors. This framework allows us to treat a broad range of connectivities on equal footing. For every case studied, we identify an optimal DMI strength for which the ergotopy dynamics exhibit an ideal sinusoidal behavior, periodically reaching the theoretical maximum $2n\hbar\omega_{0}$ and discharging fully to the passive state with negligible residual energy $\approx 0$. The optimal values of DMI interaction for different choices on $n$ and coordination number $c$ are summarized in Table 1.
To illustrate these general trends, Figure 10(d) presents results for three representative Platonic solids embedded in this framework- the tetrahedron $(n=4,c=3)$, the octahedron $(n=6,c=4)$, and the icosahedron $(n=12,c=5)$. The tetrahedral case in Figure LABEL:fig:_4_qubits_graph demonstrates that even the smallest highly symmetric geometry supports a perfect sinusoidal ergotropy cycle reaching maximum energy $2n\hbar\omega_{0}=8$ and discharging to its ground state, optimized at $D=1.96$ (Figure LABEL:4_qubits_ergotropy). The octahedron case in Figure LABEL:fig:_6_qubits_graph confirms that increasing coordination to $c=4$ stabilizes the oscillations, again yielding a full discharge at each cycle at $D=1.96$ (Figure LABEL:6_qubits_ergotropy). The maximum energy stored by this system is $2n\hbar\omega_{0}=12$. Motivated by this observation, we generalize our analysis to c-regular graphs [Kim2006, meringer1999fast] constructed using [SciPyProceedings_11], varying both the number of vertices $n$ and the degree $c$. For each graph, we tune the DMI strength $D$ to identify the value at which the ergotropy exhibits a sinusoidal behavior. The interaction Hamiltonian for such c-regular graphs is given by
$H_{int}$ $=J\sum_{(i,j)\in edges}{(1+\delta)\sigma^{x}_{i}\sigma^{x}_{j}+(1-\delta)\sigma^{y}_{i}\sigma^{y}_{j}+\Delta\sigma^{z}_{i}\sigma^{z}_{j}}$ (17)
$+D\sum_{(i,j)\in edges}{\sigma^{x}_{i}\sigma^{y}_{j}-\sigma^{y}_{i}\sigma^{x}_{j}}$
where edges denote the set of all connected node pairs. A regular graph can be generated from $n$ and $c$ provided $c<n$ and $n\times c$ is even. To construct edges, the $n$ nodes are arranged on a circle. For even $c$, each node $i$ is connected to its $\frac{c}{2}$ nearest neighbors in both clockwise and counter-clockwise directions. For odd $c$, each node is connected to its (c1)/2(c-1)/2 nearest neighbors in both directions and additionally to the node opposite it, $(i+\frac{n}{2})$.
Among all tested structures, the icosahedron (Figure LABEL:fig:_icosahedron) provides the clearest evidence of how symmetry governs efficient operation. Its five-fold coordination enables a highly uniform distribution of interactions, which translates into perfectly sinusoidal charging–discharging cycles with a complete reset to the ground state after each period. Unlike the extended cube or even the cuboctahedron, the icosahedron shows no accumulation of residual energy, ensuring that every cycle begins from the true passive state. Moreover, it reaches the maximum ergotropy value $2n\hbar\omega_{0}=24$ more rapidly than less symmetric 12-qubit structures, highlighting that carefully chosen connectivity can offset the efficiency loss normally associated with scaling. This result highlights that high geometric symmetry can recover the efficiency that is otherwise degraded in larger systems.
Taken together, the results from both $c$-regular graphs and Platonic solids demonstrate a general design principle, i.e., highly symmetric connectivities- whether represented as regular graphs or as polyhedral skeletons- preserve ideal sinusoidal ergotropy cycles across different system sizes. In particular, the icosahedron highlights that scalability to larger qubit numbers is possible without sacrificing performance, provided the underlying connectivity remains highly regular and symmetric.
IV Conclusion
@@
In this work, we studied the charging performance of Heisenberg spin chain quantum batteries (QBs) under the influence of Dzyaloshinskii–Moriya interaction (DMI). We compared three distinct spin chain configurations- open boundary, closed boundary, and supercube- under both Ising and XXZ Hamiltonians. Among these, the supercube consistently exhibited superior performance. For the Ising model, DMI enhanced ergotropy in supercube spin chain. In the XXZ model, supercube achieved a stable, near-ideal sinusoidal charging–discharging cycle, enabling periodic charging with negligible residual energy per cycle at optimal interaction strengths $D=1.7$ and $J=1$. We also examined the role of the battery Hamiltonian $H_{B}$ in the charging process through the parameter $\lambda$. For all configurations and models, increasing $\lambda$ reduced the peak ergotropy. Interestingly, in the XXZ model, ergotropy for $D=0$ and $\lambda>0$ remained higher than the $\lambda=0$ case for most of the evolution.
Beyond these baseline comparisons, we investigated structural modifications of the supercube. Adding extra interactions between qubits disrupted the clean oscillatory behavior observed in the XXZ model, preventing full discharge to the ground state and leading to residual stored energy. To address this, we systematically extended our study to $c$-regular graphs with system sizes from $3$ to $12$ qubits. Within this framework, we analyzed both generic connectivities and highly symmetric geometries, including the tetrahedron, octahedron, and icosahedron. Remarkably, all of these structures displayed ideal sinusoidal ergotropy cycles at suitable DMI strengths. In particular, the icosahedral geometry- the largest and most symmetric case studied- stood out by fully eliminating residual energy and reaching the maximum ergotropy, thereby recovering the efficiency lost in less symmetric extensions.
Although entanglement often influences QB performance, prior work has shown that entanglement generation is not strictly necessary for optimal work extraction [Hovhannisyan2013]. Consistent with this, our study did not explicitly analyze entanglement, yet demonstrated substantial enhancements in ergotropy and power without invoking it. Overall, our findings provide a roadmap for designing scalable and efficient quantum batteries. Spin chain architectures with symmetric connectivities- including the supercube, Platonic solids, and other $c$-regular graphs- achieve high-efficiency, near-ideal charging cycles without requiring entanglement, offering a promising route toward next-generation quantum energy storage devices.
@@
The used codes and data sets are available in the GitHub repository. The codes provided use the QuantumToolbox package in Python and Julia. [QuantumToolbox-jl2025, lambert2024qutip].
Appendix A Battery performance using energy level population dynamics
(a) Population dynamics of quantum battery (QB) networks. Top: Open boundary, closed boundary, and supercube spin chains with $D=0,1.7$ for (a) Ising and (b) XXZ models. Bottom: (c) Tetrahedron geometry and (d) comparison between ground and highest excited eigenstates in the XXZ model.
In this section, we examine the population dynamics of the three 8-qubit QB configurations, namely, open boundary, closed boundary, and supercube spin chains, as well as the tetrahedron QB network, to provide additional insight into the geometries discussed in Section III.2.4. Fig. LABEL:fig:ising_pop and LABEL:fig:xxz_pop shows the time evolution of population across eigenstates for the 8-qubit configurations, and Fig. LABEL:fig:tetrahedron shows the corresponding behavior for the tetrahedron network. In the 3D plots, the ground state is indexed as the 1st eigenvector and the highest excited state as the $(2^{n})^{\mathrm{th}}$ eigenvector along the x-axis.
Population inversion in QB systems is analogous to optical pumping in lasers. Recent studies show that dissipative processes can drive quantum systems to equilibrium states featuring population inversion, where the charged state can be maintained without continuous driving once equilibrium is reached [barra2019dissipative]. This behavior is especially relevant for atom–cavity QED systems, where laser-driven three-level atoms interacting with cavities in the dispersive regime achieve inversion through engineered dissipation [beleno2024laser]. Analyzing eigenstate population evolution in our QB networks therefore provides deeper understanding of charging routes and energy-storage efficiency in multi-qubit systems.
In the Ising model without DMI, excitations from the ground state are broadly distributed: all eigenstates except the ground and highest excited states become equally populated at intermediate times, reflected in the nearly uniform blue surface in Fig. LABEL:fig:ising_pop. The ground state and the highest excited state show periodic peaks, but population spreading over intermediate levels prevents smooth sinusoidal ergotropy oscillations, even though partial inversion appears. When DMI is introduced, this behavior becomes even more uniform, destroying selectivity entirely. The periodicity of peaks in the ground and highest excited states disappears, and the system no longer reaches the excited state, causing strong suppression of ergotropy oscillations as seen in Fig. 4. This behavior is consistent across all three geometries, indicating that the Ising model is insensitive to connectivity patterns.
In contrast, the XXZ model without DMI never fully excites the system to its highest energy state, as shown in Fig. LABEL:fig:xxz_pop. Open and closed chains show population spreading over many eigenstates. Remarkably, in the supercube configuration the population remains strongly concentrated in the ground state, indicating that its connectivity restricts spreading. Upon introducing DMI, open and closed chains continue to show broad population distribution with no pronounced peaks, consistent with the absence of oscillatory ergotropy. Strikingly, the supercube becomes the only configuration that exhibits clear periodic population inversion, with sharp, regularly spaced peaks alternating between the ground and highest excited states, while the intermediate eigenstates remain weakly populated. This selective excitation pathway reproduces the oscillation pattern in Fig. 7(c), highlighting the unique periodic energy-transfer mechanism supported by the supercube.
We further examine the tetrahedron QB network (a 4-qubit interaction graph) in the XXZ model to determine whether similar selective periodic inversion appears in other geometric configurations presented in Section III.2.4. As with the supercube for $D=1.7$, the tetrahedron shows clear and regular population inversion, with sharp periodic peaks corresponding to the ground and highest excited states, as seen in Fig. LABEL:fig:tetrahedron. Weakly populated intermediate eigenstates again indicate a well-defined excitation pathway. Fig. LABEL:fig:line_tetrahedron shows complete inversion via periodic peaks of the 1st and 16th eigenvectors. Similar population-inversion behavior is observed in the other graph networks listed in Table 1 and shown in Fig. 10(d).